# Model-oriented Graph Distances via Partially Ordered Sets

### v1

@@ E.3 MPDAGs
By construction, $\mathcal{G}_{1}^{\prime\prime}$ and $\mathcal{G}_{2}^{\prime\prime}$ are both covered by $\mathcal{G}_{3}^{\prime\prime}$ (see text in Section 4.4). However, there exists no graph covered by both $\mathcal{G}_{1}^{\prime\prime}$ and $\mathcal{G}_{2}^{\prime\prime}$. To see why, the only two graphs that are covered by $\mathcal{G}_{1}^{\prime\prime}$ are $\mathcal{G}_{1}^{\prime\prime\prime}$ and $\mathcal{G}_{2}^{\prime\prime\prime}$ below:
@@
Appendix F Theoretical results for CPDAGs
F.1 Lower bound on $d_{\mathcal{L}}$: proof of Property 2
Proof of Property 2.
Any path from $\mathcal{G}_{s}$ from $\mathcal{G}_{t}$ must involve removing edges that are in the skeleton of $\mathcal{G}_{s}$ and not in the skeleton of $\mathcal{G}_{t}$ and adding edges that are in the skeleton of $\mathcal{G}_{t}$ but not in the skeleton of $\mathcal{G}_{s}$. Since every edge addition or deletion increases the path length by one, we have that $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\mathrm{sk}(\mathcal{G}_{s})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G}_{s})|$.
The lower bound above does not account for v-structures that are present in one but not in the other. Specifically, suppose $\mathcal{G}_{s}$ has a v-structure $i\rightarrow j\leftarrow k$ with $i$ and $k$ not connected, and $\mathcal{G}_{t}$ does not have this v-structure, but has an edge between $i$ and $j$, an edge between $j$ and $k$, and no edge between $i$ and $k$. Thus, any path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ must remove this v-structure, which can only be done if and only if at some step at least one of the following happens: $(i)$ the edge between $i$ and $j$ and/or between $j$ and $k$ is removed, and $(ii)$ there is a covering edge that is added between $i$ and $k$. For both $(i)$ and $(j)$, an extra operation afterwards is needed: for $(i)$, it is to add back the edge, and for $(ii)$, it is to remove the covered edge. Similar argument can be made if $\mathcal{G}_{t}$ has a v-structure that $\mathcal{G}_{s}$ does not.
It is natural to define the set:
$\begin{aligned} S:=\Big\{(a,b,c)\subseteq V\times V\times V:&(a,b),(b,c)\subseteq\operatorname{sk}(\mathcal{G}_{s})\cap\operatorname{sk}(\mathcal{G}_{t}),(a,c)\not\in\operatorname{sk}(\mathcal{G}_{s})\cup\operatorname{sk}(\mathcal{G}_{t}),\\ &\text{exactly one of }(a,b,c)\in\operatorname{v-str}(\mathcal{G}_{s})\text{ and }(a,b,c)\in\operatorname{v-str}(\mathcal{G}_{t})\Big\},\end{aligned}$
and say that associated with every element in $S$, we need at least two operations. That is, in addition to $|\mathrm{sk}(\mathcal{G}_{s})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G}_{s})|$, at least $2|S(\mathcal{G}_{s},\mathcal{G}_{t})|$ additional steps must be involved in any path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$. However, $2|S|$ may overcount the number of operations needed. For example, removing an edge between $i$ and $j$ may account for elements $(i,j,k)$ and $(i,j,m)$ in $S$, or adding a covered edge between $i$ and $k$ may account for elements $(i,j,k)$ and $(i,m,k)$ in $S$.
To avoid overcounting the number of operations, we consider the following strategy. For any pair $(i,j)$, we define two operations. The first, denoted $\operatorname{OP}_{(i,j),\text{CE}}(\cdot)$, takes a subset of $S$, and removes all the elements of the form $(i,k,j)$ or $(j,k,i)$ for any $k$; this operation is associated with adding a covered edge between $i$ and $j$. The second, denoted $\operatorname{OP}_{(i,j),\text{ER}}(\cdot)$, removes all elements of the form $(i,j,k)$, $(k,j,i)$, $(j,i,k),(k,i,j)$ for any $k$; this operation is associated with deleting an edge between $i$ and $j$. Let $\mathrm{OP}=\{\operatorname{OP}_{(i,j),\text{CE}},\operatorname{OP}_{(i,j),\text{ER}}\text{ for all }\text{ vertices }(i,j)\}$ be the collection of all such operations. Our goal is to find the smallest number of operations from $\mathrm{OP}$ such that, when applied consecutively to $S$, the resulting set is empty. That is, we seek a sequence of operations $(\mathrm{op}_{1},\dots,\mathrm{op}_{m})$ where each step $\mathrm{op}_{i}\in\mathrm{OP}$, such that $\mathrm{op}_{1}\odot\dots\odot\mathrm{op}_{m}(S)=\emptyset$, and $m$ is minimized. Let the minimum be $m_{\text{op}}$. Then, it follows that:
$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\mathrm{sk}(\mathcal{G}_{s})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G}_{s})|+2m_{\text{op}}.$
F.2 Simplifications under polytree assumptions: proof of Theorem 3
The result is based on the following lemmas, which we prove first.
Lemma 3.
Let $\mathcal{G}_{a}$ and $\mathcal{G}_{b}$ be two graphs in $\mathfrak{G}$, the set of polytree CPDAGs over a common node set, with $\mathcal{G}_{a}\preceq\mathcal{G}_{b}$. Then, for every DAG $\mathcal{D}_{b}$ represented by $\mathcal{G}_{b}$, there exists a DAG $\mathcal{D}_{a}$ represented by $\mathcal{G}_{a}$, such that $\mathcal{D}_{a}$ is a directed subgraph of $\mathcal{D}_{b}$.
Proof of Lemma 3.
First, note that $\mathrm{sk}(\mathcal{G}_{a})\subseteq\mathrm{sk}(\mathcal{G}_{b})$; otherwise conditional dependencies of $\mathcal{G}_{a}$ are not a subset of those of $\mathcal{G}_{b}$. Further, if $\mathcal{G}_{a}$ has a v-structure $i\rightarrow j\leftarrow k$, then $\mathcal{G}_{b}$ must also have this v-structure. Suppose not, then, since $\mathcal{G}_{b}$ cannot have cycles and its skeleton is a superset of $\mathcal{G}_{a}$, then $\mathcal{G}_{b}$ must have an edge between $i$ to $j$ and $j$ to $k$ without an edge between $i$ to $k$. This leads to a contradiction, however, since $\mathcal{G}_{a}$ encodes the conditional dependencies $X_{i}\not\perp X_{j}|X_{S}$ for every set $S$ that contains $k$, while $\mathcal{G}_{b}$ not has have this dependency for a set $S$. Similarly, if $\mathcal{G}_{b}$ has a v-structure $i\rightarrow j\leftarrow k$, then either $\mathcal{G}_{a}$ has the same v-structure or the edge $i$ to $j$ or $j$ to $k$ is missing in $\mathcal{G}_{a}$. Suppose not, that is $i$ and $j$, and $j$ and $k$ are connected but not in a v-structure form. Then, $\mathcal{G}_{a}$ implies the conditional dependency $X_{i}\not\perp X_{j}|S$ for every $S$ that does not contain $k$. However, $\mathcal{G}_{b}$ implies the conditional independence $X_{i}\perp X_{j}|X_{S}$ for some $S$ that does not contain $k$.
Using the facts above, we conclude the desired result. Let $\mathcal{E}=\{(i,j)\}$ be collection of nodes that are connected in $\mathcal{G}_{b}$ but not in $\mathcal{G}_{a}$. By the analysis above, any v-structure in $\mathcal{G}_{a}$ is also present in $\mathcal{G}_{b}$. Furthermore, every v-structure $i\rightarrow k\leftarrow j$ in $\mathcal{G}_{b}$ is either present in $\mathcal{G}_{a}$ or at least one of the edges between pair of nodes $(i,k)$ and $(j,k)$ is missing. Consider any DAG $\mathcal{D}_{b}$ represented by $\mathcal{G}_{b}$. Let $\mathcal{D}_{a}$ be the DAG that is obtained by deleting the edges in $\mathcal{D}_{b}$ among pairs of nodes $\mathcal{E}$. By construction, $\mathcal{D}_{a}$ has the same skeleton as any DAG in $\mathcal{G}_{a}$. Since $\mathcal{D}_{a}$ is a polytree, every structure present in $\mathcal{D}_{a}$ or in $\mathcal{G}_{a}$ are also in $\mathcal{G}_{b}$. Suppose as a point of contradiction that there exists a v-structure $i\rightarrow k\leftarrow j$ that is in $\mathcal{D}_{a}$ and not in $\mathcal{G}_{a}$. Since this v-structure must also be present in $\mathcal{G}_{b}$, and from the previous statement, then one of the edges between pairs of nodes $(i,k)$ or $(j,k)$ must be missing in $\mathcal{G}_{a}$, which contradicts $\mathcal{D}_{a}$ and $\mathcal{G}_{a}$ have the same skeleton. A similar argument allows to conclude that it is not possible that $\mathcal{G}_{a}$ has a v-structure that is not present in $\mathcal{D}_{a}$. Thus, $\mathcal{D}_{a}$ and $\mathcal{G}_{a}$ have the same skeleton and v-structures. We conclude that $\mathcal{D}_{a}$ must be a DAG represented by $\mathcal{G}_{a}$. ∎
Lemma 4.
The poset $\mathcal{L}_{\text{poly}}$ underlying CPDAGs that are polytrees is graded with $\mathrm{rank}(\cdot)$ measuring the number of edges.
@@ F.2 Simplifications under polytree assumptions: proof of Theorem 3
Suppose as a point of contradiction that $\mathcal{G}_{1},\mathcal{G}_{2}\in\mathcal{G}_{\text{poly}}$ and $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$ with $\mathrm{rank}(\mathcal{G}_{1})+1<\mathrm{rank}(\mathcal{G}_{2})$. By Lemma 3, for every DAG $\mathcal{D}_{2}=(\mathrm{V},E_{2})$ represented by $\mathcal{G}_{2}$, there exists a DAG $\mathcal{D}_{1}=(\mathrm{V},E_{1})$ represented by $\mathcal{G}_{1}$ such that $E_{1}\subseteq E_{2}$. Since $\mathrm{rank}(\mathcal{G}_{1})+1<\mathrm{rank}(\mathcal{G}_{2})$, there exists $E_{3}$, satisfying $E_{1}\subset E_{3}\subset E_{2}$. Let $\mathcal{G}_{3}$ be the CPDAG obtained by completing the DAG $\mathcal{D}_{3}=(\mathrm{V},E_{3})$. By construction, $\mathcal{G}_{1}\preceq\mathcal{G}_{3}\preceq\mathcal{G}_{2}$ with $\mathcal{G}_{3}$ being distinct from $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. This leads to a contradiction.
Proof of Theorem 3.
We will show that the poset is lower semimodular and then appeal to Theorem 2 to conclude the desired result.
Let $\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3}\in\mathcal{G}_{\text{poly}}$ be three CPDAGs that are polytrees. Suppose $\mathcal{G}_{1}\lessdot\mathcal{G}_{3}$ and $\mathcal{G}_{2}\lessdot\mathcal{G}_{3}$. Consider any DAG $\mathcal{D}_{3}=(V,E_{3})$ represented by $\mathcal{G}_{3}$. By Lemma 3, there exist DAGs $\mathcal{D}_{1}=(V,E_{1})$ and $\mathcal{D}_{2}=(V,E_{2})$, represented by $\mathcal{G}_{2}$ and $\mathcal{G}_{1}$, respectively, such that $E_{1}\subseteq E_{3}$ and $E_{2}\subseteq E_{3}$. Since $\mathcal{L}_{\text{poly}}$ is graded by Lemma 4, we have that $E_{3}\setminus E_{1}=e_{1}$ and $E_{3}\setminus{E}_{2}=e_{2}$, where $|e_{1}|=|e_{2}|=1$ and $e_{1}\neq e_{2}$. Let $\tilde{\mathcal{D}}=(\mathrm{V},E_{1}\cap E_{2})$, and define $\tilde{\mathcal{G}}$ to be the CPDAG obtained by completing $\tilde{\mathcal{D}}$. By construction, $\tilde{\mathcal{G}}$ is a polytree and satisfies $\tilde{\mathcal{G}}\preceq\mathcal{G}_{1}$ and $\tilde{\mathcal{G}}\preceq\mathcal{G}_{2}$. Since $\mathrm{rank}(\tilde{\mathcal{G}})=\mathrm{rank}(\mathcal{G}_{1})-1=\mathrm{rank}(\mathcal{G}_{2})-1$, we have that $\tilde{\mathcal{G}}\lessdot\mathcal{G}_{1}$ and $\tilde{\mathcal{G}}\lessdot\mathcal{G}_{2}$. ∎
F.3 Additional theoretical results
Lemma 5.

### v2

@@ E.4 MPDAGs
By construction, $\mathcal{G}_{1}^{\prime\prime}$ and $\mathcal{G}_{2}^{\prime\prime}$ are both covered by $\mathcal{G}_{3}^{\prime\prime}$ (see Section˜4.4). However, there exists no graph covered by both $\mathcal{G}_{1}^{\prime\prime}$ and $\mathcal{G}_{2}^{\prime\prime}$. To see why, the only two graphs that are covered by $\mathcal{G}_{1}^{\prime\prime}$ are $\mathcal{G}_{1}^{\prime\prime\prime}$ and $\mathcal{G}_{2}^{\prime\prime\prime}$ below:
@@
Finally, we show that the MPDAG poset need not be graded.
Proof of Proposition˜8.
To prove that the poset $\mathcal{L}$ underlying MPDAGs is not graded, it suffices to find two maximal chains between a pair of graphs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ of unequal lengths. We first show this under $|\mathrm{V}|=4$. Fig.˜S9 presents a sequence of graphs that specifies a maximal chain from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$. This chain has length $3$. Meanwhile, Fig.˜S10 presents another maximal chain from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ that has length 5. This shows that $\mathcal{L}$ is not graded. When $|\mathrm{V}|>4$, the same argument works by adding in the remaining vertices as isolated vertices with degree zero. ∎
Figure S9: A maximal chain between MPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, where each graph covers its predecessor. To see $\mathcal{G}_{3}\lessdot\mathcal{G}_{4}$, observe that (i) $\mathcal{G}_{3}\prec\mathcal{G}_{4}$ and (ii) removing any directed edge from $\mathcal{G}_{4}$ introduces a structure that is forbidden from any MPDAG (see Fig.˜2), so no graph sits between $\mathcal{G}_{3}$ and $\mathcal{G}_{4}$.
Figure S10: A maximal chain between MPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, where each graph covers its predecessor.
E.5 Polytree MPDAGs
We first prove an auxiliary lemma.
Lemma S5.
Let $\mathcal{G}$ be a polytree MPDAG. By removing any edge from $\mathcal{G}$, we obtain an MPDAG $\mathcal{G}^{\prime}$ that satisfies $\mathcal{G}^{\prime}\prec\mathcal{G}$.
@@ E.5 Polytree MPDAGs
Using the characterization of valid MPDAGs in Section˜B.2, $\mathcal{G}^{\prime}$ is still an MPDAG: it is a PDAG, and further (i) the chordality condition is satisfied, and (ii) no forbidden structure is introduced. It then follows from Proposition˜7 that $\mathcal{G}^{\prime}\prec\mathcal{G}$. ∎
Proof of Theorem˜7.
  1. 1.
    This directly follows from definition.
  2. 2.
    Let $\mathcal{G}_{1},\mathcal{G}_{2}$ be polytree MPDAGs with $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$. By Proposition˜9, we have $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})+1\leq\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})$. For contradiction, suppose $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})+1<\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})$, or equivalently, $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})+2\leq\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})$. There are three cases to consider and we show that there is a contradiction in each case.
    • Case I
      Two or more pairs of vertices are connected in $\mathcal{G}_{2}$ but not in $\mathcal{G}_{1}$. Let $(u,v)$ be such a pair. Using Lemma˜S5, removing the edge between $u,v$ from $\mathcal{G}_{2}$ leads to an MPDAG $\tilde{\mathcal{G}}$ with $\tilde{\mathcal{G}}\prec\mathcal{G}_{2}$. It also follows from Lemma˜S5 that $\mathcal{G}_{1}\prec\tilde{\mathcal{G}}$. This contradicts $\mathcal{G}_{1}\lessdot\mathcal{G}_{2}$.
    • Case II
      There is at least one pair of vertices that is connected via an undirected edge in $\mathcal{G}_{2}$ but is connected via a directed edge in $\mathcal{G}_{1}$, and there is at least one pair of vertices, say $(u,v)$, that are connected in $\mathcal{G}_{2}$ but not in $\mathcal{G}_{1}$. By an argument similar to Case I, removing the edge between $u,v$ from $\mathcal{G}_{2}$ yields an MPDAG $\tilde{\mathcal{G}}$ satisfying $\mathcal{G}_{1}\prec\tilde{\mathcal{G}}\prec\mathcal{G}_{2}$ and hence a contradiction.
    • Case III
      There are at least two pairs of vertices, say $(i,j)$ and $(k,l)$, that are connected by an undirected edge in $\mathcal{G}_{2}$ but by a directed edge in $\mathcal{G}_{1}$. Suppose $i\to j$ and $k\to l$ appear in $\mathcal{G}_{1}$. Let $\mathcal{C}$ be the CPDAG that corresponds to $\mathcal{G}_{2}$. Because $\mathcal{G}_{2}$ is an MPDAG, it must result from imposing a set of edge orientations $\mathcal{E}$ (can be empty) on $\mathcal{C}$ and completing Meek’s rules. Let $\tilde{\mathcal{G}}$ be the MPDAG that represents $\mathcal{C}$ with orientations $\mathcal{E}\cup\{i\rightarrow j\}$. Following the argument from Case II, we have that $\mathcal{G}_{1}\preceq\tilde{\mathcal{G}}\preceq\mathcal{G}_{2}$. Due to $i\rightarrow j$ in $\tilde{\mathcal{G}}$ but iji-j in $\mathcal{G}_{2}$, we know $\tilde{\mathcal{G}}\neq\mathcal{G}_{2}$. To prove the result, we will show that $(i,j)$ and $(k,l)$ can be chosen in a way that ensures $\tilde{\mathcal{G}}\neq\mathcal{G}_{1}$.
      More concretely, we show that $(i,j)$ and $(k,l)$ can be chosen such that klk-l instead of $k\rightarrow l$ appears in $\tilde{\mathcal{G}}$. To see this, note that in a polytree, the only rule in Meek (1995) that can be applied when directing an edge is the following: if there is a structure of the form $u\rightarrow v-w$, direct the edge vwv-w as $v\rightarrow w$. By construction of $\tilde{\mathcal{G}}$, klk-l becomes $k\rightarrow l$ in $\tilde{\mathcal{G}}$ only if it is compelled by this rule. Therefore, if $k\rightarrow l$ appears in $\tilde{\mathcal{G}}$, then there must exist an undirected path $i-j-\dots-k-l$ ($j$ may be identical to $k$) in $\mathcal{G}_{2}$ such that one ore more applications of the rule propagates $i\rightarrow j$ to $k\rightarrow l$ in $\tilde{\mathcal{G}}$. However, if we were to swap the choices of $(i,j)$ and $(k,l)$, then we would have $i\rightarrow j$ and klk-l in $\tilde{\mathcal{G}}$, which finishes the proof.
  3. 3.
    Let $\mathcal{G}_{1},\mathcal{G}_{2},\mathcal{G}_{3}$ be polytree MPDAGs with $\mathcal{G}_{2},\mathcal{G}_{3}\lessdot\mathcal{G}_{1}$ and $\mathcal{G}_{2}\neq\mathcal{G}_{3}$. By the previous result, we know
    $\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{2})=\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{3})=\operatorname{\mathrm{pseudo-rank}}(\mathcal{G}_{1})-1.$
    Therefore, $\mathcal{G}_{2}$ and $\mathcal{G}_{3}$ differ from $\mathcal{G}_{1}$ by having either a single directed edge from $\mathcal{G}_{1}$ missing or having a single directed edge that is undirected in $\mathcal{G}_{1}$. Up to relabeling we therefore have three cases so let i) be the case where $\mathcal{G}_{2}$ and $\mathcal{G}_{3}$ both have an additional directed edge that is undirected in $\mathcal{G}_{1}$, ii) only $\mathcal{G}_{2}$ has such an edge and iii) neither do.
    For the first case, note that when directing an undirected edge in an MPDAG that is a polytree, the resulting graph is guaranteed to be a polytree and therefore chordal. It is also an MPDAG unless the directed edge is now part of a forbidden structure of the form $\to v$- since the other three forbidden structure cannot exist in a polytree. Since $\mathcal{G}_{2}$ and $\mathcal{G}_{3}$ are MPDAGs by assumption, it follows that the edge directed in $\mathcal{G}_{2}$ (or $\mathcal{G}_{3}$) when compared to $\mathcal{G}_{1}$ does not point into another undirected edge in $\mathcal{G}_{1}$ or $\mathcal{G}_{2}$, respectively $\mathcal{G}_{3}$. Therefore, we can direct the edge directed in $\mathcal{G}_{3}$ also in $\mathcal{G}_{2}$ to obtain $\mathcal{G}_{4}$. By construction, $\mathcal{G}_{4}\lessdot\mathcal{G}_{2},\mathcal{G}_{3}$. For the second and third case, note that we can remove any directed edge from an MPDAG that is a polytree by Lemma S5. In both Case ii) and iii), we can therefore construct $\mathcal{G}_{4}$ by deleting the directed edge missing in $\mathcal{G}_{3}$ from $\mathcal{G}_{2}$. By construction, $\mathcal{G}_{4}\lessdot\mathcal{G}_{2},\mathcal{G}_{3}$.
  4. 4.
    From Result 2, we know $\mathrm{P}_{\mathcal{L}_{\mathrm{poly}},\operatorname{\mathrm{pseudo}}}=\mathrm{P}_{\mathcal{L}_{\mathrm{poly}}}$ and hence $d_{\mathcal{L}_{\mathrm{poly}},\operatorname{\mathrm{pseudo}}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}_{\mathrm{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})$. Further, by Result 3 and Theorem˜2, we also have $d_{\mathcal{L}_{\mathrm{poly}}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L}_{\mathrm{poly}},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$.
Appendix F Further results for CPDAGs
F.1 Lower bound on $d_{\mathcal{L}}$: proof of Proposition˜6
Proof of lower bound in Proposition˜6.
Any path from $\mathcal{G}_{s}$ from $\mathcal{G}_{t}$ must involve removing edges that are in the skeleton of $\mathcal{G}_{s}$ and not in the skeleton of $\mathcal{G}_{t}$ and adding edges that are in the skeleton of $\mathcal{G}_{t}$ but not in the skeleton of $\mathcal{G}_{s}$. Since every edge addition or deletion increases the path length by one, we have that $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\operatorname{sk}(\mathcal{G}_{s})\triangle\operatorname{sk}(\mathcal{G}_{t})|$.
The lower bound above does not account for v-structures that are present in one but not in the other. Specifically, suppose $\mathcal{G}_{s}$ has a v-structure $i\rightarrow j\leftarrow k$ with $i$ and $k$ not connected, and $\mathcal{G}_{t}$ does not have this v-structure, but has an edge between $i$ and $j$, an edge between $j$ and $k$, and no edge between $i$ and $k$. Thus, any path from $\mathcal{G}_{s}$ to $\mathcal{G}_{t}$ must remove this v-structure, which can only be done if and only if at some step at least one of the following happens: (i) the edge between $i$ and $j$ and/or between $j$ and $k$ is removed (this is the operation $\operatorname{OP}_{(i,j),\text{ER}}$); or (ii) there is a covering edge that is added between $i$ and $k$ (this is the operation $\operatorname{OP}_{(i,j),\text{CE}}$). For both (i) and (ii), an extra operation afterward is needed: for (i), it is to add back the edge; and for (ii), it is to remove the covered edge. Similar argument can be made if $\mathcal{G}_{t}$ has a v-structure that $\mathcal{G}_{s}$ does not.
The set $\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})$ precisely defines the set of relevant triplets described in the previous paragraph. Let $\mathrm{OP}:=\{\operatorname{OP}_{(i,j),\text{CE}},\operatorname{OP}_{(i,j),\text{ER}}\forall i,j\in\mathrm{V}\}$. Our goal then becomes finding the smallest number of operations, from the set of operations $\mathrm{OP}$ such that, when applied consecutively to $S(\mathcal{G}_{s},\mathcal{G}_{t})$, the resulting set is empty. In other words, we obtain that:
$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\mathrm{sk}(\mathcal{G}_{s})\setminus\mathrm{sk}(\mathcal{G}_{t})|+|\mathrm{sk}(\mathcal{G}_{t})\setminus\mathrm{sk}(\mathcal{G}_{s})|+2m_{\text{opt}}(\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})).$
where
$m_{\mathrm{op}}(\mathrm{S}):=\min\{m\in\mathbb{N}:\exists\mathrm{op}_{1},\dots,\mathrm{op}_{m}\text{ such that }\mathrm{op}_{1}\circ\dots\circ\mathrm{op}_{m}(\mathrm{S})=\emptyset\}.$
Consider the partition $\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})=\mathrm{S}_{1}\cup\dots\mathrm{S}_{L}$. Since triplets in distinct partitions do not share any common vertices, we have that:
$m_{\mathrm{op}}(\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t}))=\sum_{\ell=1}^{L}m_{\mathrm{opt}}(\mathrm{S}_{\ell}).$
F.2 Additional results
Lemma S6.